BRVK operational status:
IDA station BRVK has been decommissioned, last data was day 2019,235-05:34. IDA engineers have installed new station BORK nearby. Details to be posted to BORK pages soon.
The station operator replaced the LAN extenders and ethernet surge arrestors, restoring data acquistion on day 285-06:00.
Power was restored day 2018,270-05:01, however there are still communication problems between the recording room and vault. Replacement LAN extenders and ethernet surge arrestors are being prepared to ship to the station.
The station is down as of day 2018,257-09:50 due to power problems. IDA engineers are working with the station operator to resolve this issue.
Station telemetry was down day 2017,159 (June 8) due to damages to station equipment caused by lightning (unfortunately following the work at the station). Telemetry was restored after the station operator reset the Black box(short haul modems) for the LAN extenders. IDA engineers are working with the station operator to replace the damaged equipment.
An IDA engineer visited the station to upgrade the power system and install a Setra barometer.
Telemetry and data acquisition restored with no data loss, after an IDA engineer configured a backup computer at the station after the primary station computer failed.
The station operator reports a problem with the Internet connection to the station. The internet service provider is working to resolve this problem.
An IDA engineer is currently on-site upgrading the station to a Quanterra Q330HR system. Seismometers include STS-1 and STS-2 seismometers, and a strong motion sensor.
The station operator reports moving all the station recording equipment to the bunker where the vault is located, and installed new backup batteries for the power system.
The horizontal channels from the STS-1 sensor are showing excessive spikes as of about day 2008,001 for reasons yet to be determined.
An IDA engineer visited the station and reinstalled an STS-1 seismometer in place of the STS-2 that was temporarily installed.
The East-West broadband channels are nonseismic as of about day 137-12:20 for reasons yet to be determined.
The broadband channels (BB*) are restored after the station operator replaced the STS-1 sensor with an STS-2 sensor. The STS-2 is a temporary installation until the STS-1 is fixed and sent back to the station.
Station is up after an IDA engineer replaced the ISP, restoring communication between the DAS and ISP. There are problems with the STS-1 sensors and they have been sent back to UCSD for troubleshooting/repair.
Station is down as of 02/19/05 due to communication problems between the DAS and ISP.
Station is back up after the station operator switched the uplink port on the GPS clock.
Station is down due to problems with the DAS (01/31/2005).
Station back up after an IDA engineer rebooted the DAS.
Station is down due to communication problems between the ISP and DAS.
Station operating normally.
Station operator reports problems with Power Supply. The station remains operational, however, there is no power to the GPS clock and microbarometer.
Telemetry temporarily down from 2003.07.09 through 2003.07.11 following storm. Station remained operational and NRTS data for this time period was acquired when telemetry restored.
Operator replaced DSP board using spare shipped from San Diego. Data collection of S**/LN*/EN*/VM* channels restored.
Data streams from 16 bit channels (LN*, EN*, S**, VM*) are non-seismic. Failure of DAS DSP board suspected.
Station is back up following repair of power system.
New EPROMs installed to correct DSP problem.
Station is down due to power supply problem.
Station operating normally.